# MovieLLM: Enhancing Long Video Understanding with AI-Generated Movies

### v1

@@ 4.2 Evaluation Metrics
We primarily evaluate our approach in two main aspects: the quality of generated video key frames and video understanding.
@@
We consider these two metrics to evaluate the generated results: 1) key frame consistency, which is calculated by average cosine similarity between adjacent embeddings of key frames in CLIP image space. 2) text-image alignment, which is calculated by the average cosine similarity between each key frame description and key frame itself. Besides, we use the non-reference image quality assessment metric BRISQUE score (Mittal et al., 2012) to evaluate the image quality.
@@
For video understanding, we conduct evaluations separately for short and long videos. Given that there is no benchmark specifically tailored for movie-level video understanding in existing works, we draw inspiration from the design of metrics for short video understanding and meticulously devise our long video understanding benchmark.
For short video understanding, we evaluate models on two benchmarks: a) Video-based Generative Performance Benchmarking, which includes five aspects to estimate model performance. b) Zero-shot Question-Answer Evaluation, which is a comprehensive quantitative evaluation using several commonly used open-ended QA datasets, including MSVD-QA (Chen and Dolan, 2011) and MSRVTT-QA (Xu et al., 2016).
For long video understanding, we present a benchmark tailored to this task. Our test dataset is derived from approximately 100 movies within the MovieNet database, and associated question-answer pairs are generated from both GPT-4 and human efforts. Inspired by Maaz et al. (2023), we redesign the evaluation pipeline using GPT-4. The evaluation prompt can be found in our appendix B.2. This pipeline assesses various capabilities of the model by comparing different predicted answers from different models based on the ground truth and scoring the predicted answers on a scale of 1-5 in the following three aspects:
Figure 5: Comparison of the key frames generated by ours and other methods. Our method generates rather more consistent and high-quality frames.
@@ 4.3 Main Results
We first compare the quality of key frames generated using our pipeline with similar existing methods. Then, we evaluate the performance on short video understanding and long video understanding.
@@
We compare our method based on immobilized style embedding with two similar existing methods: 1) Custom diffusion (Kumari et al., 2023), a representative multi-concept customization method. 2) Cones-V2 (Liu et al., 2023b), a multi-subject customizable image synthesis method utilizing layouts. Table. 2 that our method outperforms the existing methods across all three metrics, indicating our method generates consistent and high-quality key frames.
Short Video Understanding
The model trained on our datasets gains significant performance improvement over baseline as listed in Table. 3 and Table. 4. The results indicate that although our data is focused on long video instruction tuning, it can still enhance the model’s understanding of short videos, demonstrating the robustness of our data.
Long Video Understanding
Results on Long Video Understanding are demonstrated in Table. 5. Our method outperforms the baseline by a significant margin in aspects of overview, plot, and temporal understanding. This showcases the richness of our dataset, proving their significant effectiveness in enhancing the model’s understanding in long videos across various aspects.
@@ 4.4 Qualitative Results
Fig. 5 visualizes qualitative results on the quality of generated key frames. Our method generates rather more consistent and high-quality frames. Fig. 6 visualizes video understanding results of the baseline model and ours. Our method demonstrates a more reasonable and detailed answer in a movie understanding.
5 Related Work
Vision Language Models
With the achievements of large language models (LLMs) such as GPT-4 (OpenAI, 2023) along with their open-source alternatives like LLaMA (Touvron et al., 2023), researchers focus on leveraging the advanced language abilities of LLMs and developing the vision language models (VLMs) that integrate vision models with LLMs for cross-modality understanding. Previous representative VLMs, miniGPT-4 (Zhu et al., 2023) and LLaVA (Liu et al., 2023a), have shown great capabilities in visual chat by constructing high-quality image-instruction pairs to align the image and text dimensions. Further, VLMs are developed for video understanding (Jin et al., 2024; Li et al., 2023b; Lin et al., 2023; Bai et al., 2024; Ma et al., 2023). Video-LLaMA (Zhang et al., 2023) utilizes BLIP-2 (Li et al., 2023a) to extract video embedding for each frame and fuse video embedding of all frames by video Q-Former. Video-ChatGPT (Maaz et al., 2023) designs to extract video embedding by averaging frame-level features across temporal and spatial dimensions respectively. However, previous works require generating a large number of tokens for each frame for representation. For hour-long videos, these VLMs are unable to process extensive video sequences. To tackle this problem, LLaMA-VID (Li et al., 2023c) is proposed for long video understanding by encoding each frame with only two tokens.
Video Instruction Tuning Datasets
Preparing video-text pairs to construct video instruction tuning datasets is pivotal in facilitating the training of VLMs for video understanding. Most existing methods (Maaz et al., 2023; Tang et al., 2023; Zhong et al., 2022; Gao et al., 2021; Castro et al., 2022; Yang et al., 2023) construct a video instruction tuning dataset based on benchmark datasets (Han et al., 2023), utilizing rich annotation information. The Video-ChatGPT, for example, constructs VideoInstruct100K sourced from the ActivityNet dataset (Caba Heilbron et al., 2015). However, most existing methods lack long video data in the constructed instruction tuning datasets, resulting in poor long-video understanding capabilities of the trained models. The recently proposed LLaMA-VID addresses this by extracting data from the MovieNet dataset (Huang et al., 2020) to build an instruction dataset containing movie-level long videos. However, such datasets are limited by the types of annotations, leading to insufficient diversity in the data. Leveraging our novel pipeline, we propose a diverse movie-level video instruction tuning dataset for enhancing long video understanding.
@@ 6 Conclusion
In this paper, we propose an effective method for generating key frames of movie-level videos. Our generation process consists of three main steps: movie plot generation, style immobilization process, and video instruction data generation. The proposed data generation workflow significantly eases the challenge of producing movie-level video data for models, enhancing both the control and diversity of the generated content. Experiments conducted on existing datasets, as well as on our newly generated dataset, validate the effectiveness of our approach. Thanks to the richness of our dataset, models demonstrate capabilities previously unattainable with training on earlier datasets. We believe our data generation pipeline will make a significant contribution to the advancement of multimodal models’ understanding of long videos. Furthermore, it paves the way for the adoption of similar data generation processes across other domains.
@@ 7 Limitations
The forgetting issue associated with large language models may lead to inconsistencies in the frame descriptions generated during the first stage of our pipeline, subsequently causing discontinuities in the video scenes. To address this problem, we plan to refine the text generation component within our pipeline.
@@ Appendix A Experiment Details
A.1 Training Details
During the model training phase, we employed the original LLaMA-VID (Li et al., 2023c) configuration as the foundation for our training process. We utilized 4 NVIDIA A100 GPUs. To conserve GPU memory, we employed deepspeed with zero3 during model training, disabling tf32 and opting for fp16. The remaining parameters, including epochs and learning rates, were set according to the original LLaMA-VID configuration. For specific parameter details, please refer to the the original paper (Li et al., 2023c).
@@ Appendix C More Results of our Generated Data
We illustrate more results of our generated data in Fig 12.
@@ Appendix D More Comparison Results of Long Video Understanding
Figure 12: Examples of generated long video instruction data We use GPT-4 and guided text-to-image generation models to generate consistent key frames of move-level video with reasonable lines and corresponding question-answer pairs. These data are used to train multimodal large language models on video understanding.

### v2

@@ 4.2 Evaluation Metrics
We consider these three aspects to evaluate the generated key frames: 1) key frame consistency, which is calculated by average cosine similarity between adjacent embeddings of key frames in CLIP image space. 2) text-image alignment, which is calculated by the average cosine similarity between each key frame description and key frame itself. 2) quality of image, which is calculated by non-reference image quality assessment metric BRISQUE score (Mittal et al., 2012) to evaluate the image quality.
@@
We follow the quantitative evaluation method from Video-ChatGPT (Maaz et al., 2023). The method basically includes two types of evaluations: i) Video-based Generative Performance Benchmarking and ii) Zero-Shot Question-Answer Evaluation. The essence of these two evaluations is to employ GPT assistance to assess the model’s capabilities. Specifically, evaluations measure the accuracy of the model’s generated predictions and assigns a relative score on a scale of 1-5. Besides, in ablation studies, we also report several standard machine translation metrics including BLEU-4 (Papineni et al., 2002), METEOR (Lavie and Agarwal, 2007), ROUGE-L (f-score) (Lin, 2004). Notably, in this work, we utilize GPT-based methods for most evaluations since metrics like BLEU-4 rely heavily on surface-level n-gram matching and literal similarity, lacking deep semantic understanding and context awareness. This limits their effectiveness in recognizing synonyms, metaphors, and context-dependent meanings.
@@ 4.3 Main Results
We compare our method based on immobilized style embedding with two similar existing methods: 1) Custom diffusion (Kumari et al., 2023), a representative multi-concept customization method. 2) Cones-V2 (Liu et al., 2023b), a multi-subject customizable image synthesis method utilizing layouts. Table. 2 shows that our method outperforms the existing methods across all three metrics, indicating that our method generates more consistent and text-aligned key frames with higher image quality. Fig. 4 visualizes qualitative results on the quality of generated key frames.
Video Understanding
Table. 3 presents a comparative assessment of our approach against several leading methods on three zero-shot video question answering benchmarks. Utilizing a substantial amount of synthetic tuning data, MovieLLM shows superior performance on almost all the metrics and come close on MSVD-QA. This indicates the effectiveness of our approach as a synthetic data generation framework for instruction tuning. In Table. 4, we report results on video-based generative performance benchmark (Maaz et al., 2023). MovieLLM attain the highest performance across all evaluation metrics, outperforming previous approaches by a large margin. This futher validates the effectiveness of our method.
4.4 Ablation Studies
We conduct ablation studies to investigate mainly the impact of varying proportions of our dataset on model performance and the efficacy of our dataset compared with MovieNet. For ablation study on propotion, we tune LLaMA-VID with our data of different ratio. Table. 5 shows the results on the impact of varying proportions of our dataset on model performance. With the increase of data proportions, model performance improves accordingly. This indicates the effectiveness of our generated data. For the efficacy experiment, we construct a counterpart dataset from MovieNet as mentioned in Section 4.1 (named as part 2). Part 2 consists of the same amount of QA pairs with ours. Then we use dataset from these two sources to tune LLaMA-VID. The results are shown in Table. 6 and Table. 7. The outcomes indicate that tuning with our dataset brings an obvious performance upgradation. We find that the generation process based on consistent key frame descriptions actually leads to QA pairs that are more closely related to keyframes which improve the tuning of the model. Table. 8 shows the results on Ego4D with the standard machine translation metrics. The outcome is consistent with the previous ones. Besides, there is currently no benchmark for movie-level video understanding. Therefore, we construct one based on the real movie as mentioned. The results are provided in Table 9 which shows tuning with our dataset indeed brings higher performance on long video undersatnding.
@@ 4.5 Qualitative Results
Fig. 4 visualizes qualitative results on the quality of generated key frames. Our method generates rather more consistent and high-quality frames. For more qualitative results, please refer to our appendix.
@@ 5 Conclusion
In this paper, we propose an effective method for generating key frames of movie-level videos. Our generation process consists of three main steps: movie plot generation, style immobilization process, and video instruction data generation. The proposed data generation workflow significantly eases the challenge of producing long video data for models, enhancing both the control and diversity of the generated content. Extensive experiments are conducted to validate the effectiveness of our approach. Thanks to the richness of our dataset, models demonstrate capabilities previously unattainable with training on earlier datasets. We believe our data generation pipeline will make a significant contribution to the advancement of multimodal models’ understanding of long videos. Furthermore, it paves the way for the adoption of similar data generation processes across other domains.
@@ 6 Limitations
The memory length issue associated with large language models may lead to inconsistencies in the frame descriptions generated during the first stage of our pipeline, subsequently causing discontinuities in the video scenes. To address this problem, we plan to refine the text generation component within our pipeline. Besides, although low fps video datasets have already been effective in enhancing the capabilities of our model, we still plan to explore efficient methods for generating high fps long videos to further improve the quality of our data.
@@ Appendix A Experiment Details
A.1 Model Architecture and Training Details
Model Architecture
Here, we introduce the fundamental modules used in LLaMA-VID. LLaMA-VID basically consists of a visual encoder, a text decoder, a projector and a LLM. For visual encoder, LLaMA-VID use EVA-G (Fang et al., 2023) as its ViT-based backbone and the patch size is set to 14. For text decoder, Qformer-7b (Dai et al., 2023) is used in our experiment. For projector, one-layer MLP is used to transform the embedding into context token. For LLM, we use vicuna-7b.
Training Details
As mentioned, we conduct experiments on the third stage of LLaMA-VID. During the model training phase, we employed the original LLaMA-VID (Li et al., 2023d) configuration as the foundation for our training process. We utilized 4 NVIDIA A100 GPUs. To conserve GPU memory, we employed deepspeed with zero3 during model training, disabling tf32 and opting for fp16. The remaining parameters are shown in Table.10
Settings Stage 3
Batch size 4
Learning rate 2e-5
Learning schedule Cosine decay
Warmup ratio 0.03
Weight decay 0
Epoch 1
Optimizer AdamW
Vision encoder Freeze
Text decoder Freeze
Max token 65536
Table 10: Training settings.
@@ Appendix C More Results of our Generated Data
We illustrate more results of our generated data in Fig 5 and Fig 6.